OpenStudy (anonymous):

Annuity Problem. A house cost $250,000 cash. A purchaser will pay $50,000 cash and a sequence of 8 annual payments, the first due at the end of 3 years. If moiney is worth 7%. Find the annual payment.

OpenStudy (caominhim):

\[A = P(1-\frac{r}{100})^{n}\]

OpenStudy (caominhim):

and if the buyer pays 50,000 cash you will have 0 interest and a very happy seller

OpenStudy (anonymous):

then what's next?

OpenStudy (caominhim):

did you mean 250,000 cost? or 5,000 down?

OpenStudy (anonymous):

250, 000 cost

OpenStudy (anonymous):

how to solve it...

OpenStudy (wolf1728):

Is this really an annuity problem? Seems more like a loan or a mortgage problem.

OpenStudy (anonymous):

i need help to solve this.

OpenStudy (anonymous):

i do not know how to solve this, can someone help?

OpenStudy (tkhunny):

Basic Principles i = 0.07 v = 1/(1+i) d = iv 250000 = 50000 + Pv^3 + Pv^4 + ... + Pv^10 \(200000 = P\dfrac{v^{3} - v^{11}}{1-v} = P\dfrac{v^{3}}{d}\left(1-v^{8}\right) = P\dfrac{v^{2}}{i}\left(1-v^{8}\right)\) These never have to be tricky. Just build them.

OpenStudy (anonymous):

what is P?

OpenStudy (tkhunny):

"Find the annual payment." P = The Annual Payment Normally, I would define that up front. Must have been typing in my sleep. You must do the arithmetic. You must first understand how it came to be.